OpenStudy (anonymous):

rashell id baking a cake but does nor know if she wants to bake a round or square cake.which cake has a larger area, 9 round 8 inch cake or a sqaure 9 inch cake?how much larger pleaseee helppp.

OpenStudy (anonymous):

is the round or square bigger and a number of how much bigger

OpenStudy (anonymous):

i wanted to break it down tha was alot to read

OpenStudy (anonymous):

@yunus u have an idea

OpenStudy (anonymous):

im trying to understand what does mean "9 round 8 inch cake "

OpenStudy (anonymous):

oh im sorry let me redo that for u

OpenStudy (anonymous):

which cake is larger area a round 8 inch cake or a square 9 inch cake

OpenStudy (anonymous):

lol sorry and how mich bigger

OpenStudy (anonymous):

@yunus

OpenStudy (anonymous):

let me give my answer choices

OpenStudy (mathstudent55):

Find the area of each. You need the area of a circle and the area of a square. Do you know how to get thoses areas?

OpenStudy (anonymous):

the area of the ROUND cake is larger by about 31 in to the second power the area of the SQAUREcake is larger about 31 in to the second power the area of the ROUND cake is larger by 11 in to the second power the area of the SQAURE cake is about 11 in to the second power

OpenStudy (anonymous):

ok there si my answer choices ther please help

OpenStudy (mathstudent55):

The area of a circle is A = (pi)r^2, where r = length of radius, and pi = 3.14 (approx) The area of a square is A = s^2, where s = length of side

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

A round 8-inch cake is a cake baked in a pan that has a diameter of 8 inches. The diameter is 2 times the radius. So the round pan has a radius of 4 inches.

OpenStudy (anonymous):

ok @yunus

OpenStudy (mathstudent55):

The area of a circle with a radius of 4 inches is: A = (pi)(4^2) = 16(pi) = 50 sq in. (approx)

OpenStudy (anonymous):

okkk

OpenStudy (mathstudent55):

The area of a square with a side of 9 inches is: A = s^2 = 9^2 = 81 sq in.

OpenStudy (mathstudent55):

Now you need to comapre the round pan with 50 in^2 area to the square pan with 81 in^2 area and choose the correct answer choice.

OpenStudy (anonymous):

okkk

OpenStudy (anonymous):

@yunus what do u htink

OpenStudy (anonymous):

after this ione i will post aniother one up

OpenStudy (anonymous):

mathstudent55 is right. he found the two values of area already. :)

OpenStudy (mathstudent55):

I gave you both areas. Read the choices and see which one makes sense. Which one has the larger area, the square or the circle?

OpenStudy (anonymous):

yes he is or girl sorry

OpenStudy (anonymous):

the square

OpenStudy (anonymous):

but the question s the numbers after that

OpenStudy (mathstudent55):

Right. How much bigger is the area of the square than the area of the circle?

OpenStudy (anonymous):

ohh i think it is he. mathstudent are you male or famale?

OpenStudy (anonymous):

like the 31 or 11

OpenStudy (anonymous):

i was thikn 31

OpenStudy (mathstudent55):

There is nothing to guess. One area is 81 the other is 50. How much bigger is 81 than 50?

OpenStudy (mathstudent55):

BTW, you were thinking right

OpenStudy (anonymous):

31

OpenStudy (mathstudent55):

You'va answered it. The square is bigger by 31. Now pick the corect choice.

OpenStudy (anonymous):

is b

OpenStudy (anonymous):

thank you for al the help>:)

OpenStudy (mathstudent55):

correct